LINE SHIFT AND LINE BROADENING IN FUNDAMENTAL BANDS OF DIFFERENT ISOTOPOMERES OF CARBON MONOXIDE

\underline{S. VOIGT}, J. P. BURROWS, J. ORPHAL, Institute of Environmental Physics, University of Bremen, P. O. Box 330440, 28334 Bremen, Germany; A. KISSEL, H.-D. KRONFELDT, O. KURTZ, I. MEUSEL, and B. SUMPF, Institute of Optics, Technical University Berlin, Sekr. PN 0-1, Hardenbergstr. 36, 10623 Berlin, Germany.

In this paper we present results concerning line shift and line broadening in the fundamental band of different isotopomeres of carbon monoxide near 2058 cm^{−1}. The study includes lines of ^{12}C^{16}O and for the first time lines of ^{13}C^{16}O, ^{12}C^{18}O, 13C^{18}O, and ^{13}C^{17}O perturbed by N_{2}, O_{2}, H_{2}, D_{2}, He, Ne, Ar, Kr, and Xe. The results for Kr, H_{2}, and D_{2} represent the first experimental data for the (1←0) band of CO. For all perturbers except helium all line-shift coefficients are negative. Within the experimental uncertainties the data show no significant dependence on the isotopomere. As known for ^{12}C^{16}O for collisions with noble gases the absolute value of the line-shift coefficients increase with increasing mass of the perturber. Lines from R-branches show smaller shift coefficients compared to P-branch lines. Our absolute values for the line-shift coefficients for the (1←0) band confirm the trend that the effect for this band is smaller compared to the (2←0) band as known from former experimental data by Bouanich et al.^{a}. The nitrogen broadening will be compared to previous measurements for ^{13}C^{16}O by Voigt et al.^{b}.

aJ.-P. Bouanich, D. Bermejo, J. L. Domenech, R. Z. Martinez, J. Mol. Spectrosc. 179, 22 (1996)

bS. Voigt, S. Dreher, J. Orphal, J. P. Burrows, J. Mol. Spectrosc. 180, 359 (1996)